What is the wavelength of the X-ray emitted from a Tungsten target (Z=74) when an electron moves from the M-shell (n=3) to the K-shell (n=1)
1
Expert's answer
2020-06-08T10:13:17-0400

Given:

Z=74

n1=1

n2=2


The energy is


"\Delta E=hcR_\infty Z^2\bigg(\frac{1}{n_1^2}-\frac{1}{n_2^2}\bigg)"

The wavelength is


"\lambda=\frac{hc}{\Delta E}=\frac{1}{R_\infty Z^2\bigg(\frac{1}{n_2^2}-\frac{1}{n_1^2}\bigg)}=444\cdot10^{-6}\text{ m}."

In this equation, "R_\infty" is the  Rydberg constant, "R_\infty=10973731.57 \text{ m}^{-1}".